OpenStudy (anonymous):

What is a possible value for the missing term of the geometric sequence? 1250, ___, 50, ...

OpenStudy (mathmale):

Here, the first term of the sequence is 1250. The next term is obtained by multiplying that by the common ratio, r: 2nd term = 1250 (r) = x. 3rd term = 1250(r)(r)=50. Find r^2. Then find r. then use this r to compute the 2nd term.

OpenStudy (anonymous):

1250/50 = 25 \[\sqrt{25} = 5 \] 1250, 250, 50, 10 Thank you!

OpenStudy (mathmale):

You're welcome! But what is the value of r? (Hint: r is not a whole number; it's a fraction.)